A 0.3749 g soda ash sample is analyzed by titrating sodium carbonate with the standard 0.2388M HCl solution, requiring 49.38ml. The reaction is:

          CO32-  + 2H    →   H2O +CO2

1
Expert's answer
2021-03-24T01:11:04-0400

soda ash = sodium carbonate = Na2CO3


Solution:

The reaction is:

CO32- + 2H+ → H2O + CO2

According to the reaction above: n(CO32-) = n(H+)/2


HCl → H+ + Cl-

n(H+) = n(HCl) = CM(HCl) × V(HCl) = (0.2388 M) × (0.04938 L) = 0.011792 mol


n(CO32-) = n(H+) / 2 = 0.011792 mol / 2 = 0.005896 mol


Na2CO3 → 2Na+ + CO32-

n(Na2CO3) = n(CO32-) = 0.005896 mol


Mass of Na2CO3 = Moles of Na2CO3 × Molar mass of Na2CO3

The molar mass of Na2CO3 is 105.9888 g mol-1.

Hence,

Mass of Na2CO3 = (0.005896 mol) × (105.9888 g mol-1) = 0.6249 g

Mass of Na2CO3 = 0.6249 g


%Na2CO3 = (Mass of Na2CO3 / Mass of soda ash sample) × 100%

However,

Mass of Na2CO3 > Mass of soda ash sample

0.6249 g > 0.3749 g

Perhaps there was an error in the problem statement!

Since the percent sodium carbonate in the sample cannot be more than 100%.